OpenStudy (anonymous):

A wooden garden tray with a base area of xy square inches is filled with soil up to a certain height. The tray contains (4x2y + 6xy2 + xy) cubic inches of soil. The height of the tray is represented by the expression 4x^2y+6xy^2+xy/xy Which expression below is the correct simplified form of the height of the tray? 10xy 10xy + 1 4x + 6y + 1 4x2y + 6xy2 + 1

OpenStudy (anonymous):

@jazy

OpenStudy (anonymous):

@Callisto

OpenStudy (callisto):

\[\frac{4x^2y+6xy^2+xy}{xy} = \frac{xy(4x+6y+1)}{xy} =?\]

OpenStudy (callisto):

Actually, the question is simply: simplify the expression: 4x^2y+6xy^2+xy/xy

OpenStudy (turingtest):

yeah totally misunderstood it :S

OpenStudy (anonymous):

thanks

OpenStudy (callisto):

Welcome!~